3. Compute the moisture content of steam at 20 kPa, entropy of 6.3 kJ/kg-K.


1
Expert's answer
2022-03-09T23:38:02-0500

P=20kPa

s=6.3kJ/kgK

"s=s_f+xs_{fg}"

6.3=0.832+x7.075

6.3-0.832=x7.075

5.468=x7.075

"x=\frac{5.468}{7.075}"

x=0.77286